OpenStudy (anonymous):

Determine the restricted values for the variable when solving the equation shown here. 5/6x-3/2x=-3

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6x,2x \ne0 }\)

Parth (parthkohli):

Solve both of them to find the restricted value(s)

OpenStudy (anonymous):

@ParthKohli how do i solve those?

Parth (parthkohli):

\( \color{Black}{\Rightarrow 6x \ne 0 }\) \( \color{Black}{\Rightarrow 2x \ne 0 }\) Two different equations. I've put \(\ne\) 0, because the denominator cannot be 0.

OpenStudy (anonymous):

5/6x-3/2x=-3 5-9/6x=-3 -4/6x=-3 -2/3x=-3 -2/-3=3x 2/3=3x x=2/9

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

Parth too showed the same but I done it with step by step but he deserves a medel too

OpenStudy (anonymous):

can u give out 2 medals?